OpenStudy (anonymous):

Pyramid A is a square pyramid with a base side length of 14 inches and a height of 6 inches. Pyramid B has a volume of 3136 cubic inches. How many times bigger is the volume of pyramid B than pyramid A? What is the percent change from pyramid A to pyramid B? Provide an explanation and proof for your answer

OpenStudy (nirmalnema):

volume of sq pyramid=1/3 b^2*h

OpenStudy (nirmalnema):

volume of pyramid A = 1/3 * 196 * 6 =196*2 =392

OpenStudy (nirmalnema):

Pyramid B has a volume of 3136 cubic inches vol...a/b =3136/392=8times

OpenStudy (nirmalnema):

percent change = 392/3136 * 100 =12.5%

OpenStudy (nirmalnema):

@lowcard2 do u get it

OpenStudy (campbell_st):

well if pyramid B is 8 times larger than pyramid A... when comparing volumes, the change can't be an increase of 12.5%

OpenStudy (anonymous):

yeah is that all i right then??

OpenStudy (nirmalnema):

yup.

OpenStudy (anonymous):

thxx

OpenStudy (nirmalnema):

welcome....

OpenStudy (campbell_st):

the scaling factor is correct for the volumes... but the percentage increase isn't correct... New volume = 392 + 392 * 12.5/100 = 441 so if you say the volume of pyramid A has increased by 12.5% then the new volume is 441 cubic inches